OpenStudy (anonymous):

a warning system installation consists of two independent alarms having probabilities of operating in an emergency of 0.95 and 0.90 respectively. find the probability that at least one alarm operates in an emergency?

OpenStudy (anonymous):

2 Alarms: p(working) .95 and .90 P(at least one working) = 1 - P(neither working) = 1 - (.05)(.10) = .995

OpenStudy (anonymous):

thanks @awsomekrishn :)

OpenStudy (anonymous):

no problem :)